Height:165cm
The walls should be 7/8 of the height
The pole should be 3/4 of the height
The distance between the poles should be 5/4 the height of the height
The middle poles should have 2/3 their height between the height
The length of poles should be 1 9/11 of the width of the field
What is the area of the field
What is the area of the surrounding walls
1
Expert's answer
2019-06-14T11:11:28-0400

The height of the walls is 165*7/8 = 144.375. Since the Distance between the poles must be 5/4 of the height of the height of the Middle poles must have 2/3 of their height between the height Of the length of the poles must be 19/11 of the width of the field, then the Length of the field is multiplied by the width (19/11)*(5/4)*165 = 356.25, and the width is (5/4)*165 =206.25. Then the area of the base is equal to 356.25*206.25 =73476.5625, and the area of the walls is equal to 144.375*206.25*2+144.375*356.25*2=162421.875